# Whitehead’s problem and condensed mathematics

### v2

@@ 10. Martin’s Axiom and $\mathsf{CondAb}_{\omega_{1}}$
For a given abelian group $A$, if $\underline{A}$ is Whitehead in $\mathsf{CondAb}_{\aleph_{1}}$, then $A$ is Whitehead in the classical sense, since
@@
Let $\vec{p}=\langle p_{\eta}\mid\eta<\omega_{1}\rangle$ be a sequence of conditions in $\mathbb{P}$. By (the proof of) [5, Lemma 7.5], by thinning out $\vec{p}$ to some cofinal subsequence if necessary, we can assume that there is a free subgroup $A^{\prime}$ of $A$ that is pure in $A$ and such that $A_{p_{\eta}}\subseteq A^{\prime}$ for all $\eta<\omega_{1}$ (this is the only use of the full power of Chase’s condition in the proof). Let $B^{\prime}$ be a basis for $A^{\prime}$. Using Claim 10.5 and extending each $p_{\eta}$ if necessary, we can assume that, for all $\eta<\omega_{1}$, $A_{p_{\eta}}$ is freely generated by some finite subset $B^{\prime}_{\eta}$ of $B^{\prime}$. By thinning out $\vec{p}$ again if necessary, we can assume that the sets $\{B^{\prime}_{\eta}\mid\eta<\omega_{1}\}$ form a $\Delta$-system, with root $R$.
@@ 11. A different vantage point
To see the fifth equality, apply Lemma 3.2; for the third and fourth, respectively, apply [13, Prop. 3.5 and Prop. 3.2]. The first equality follows from [20, p. 13 (iii) (cf. p. 25)]. This leaves us only with the second. Here the idea, evidently, is to apply the Hom-tensor adjunction at the level of $\mathrm{Ext}^{1}_{\mathsf{CondAb}}$, and we may do so by the naturality of the adjunction at the $\mathrm{Hom}_{\mathsf{CondAb}}$-level, the flatness of $\mathbb{Z}[\underline{S}]$ [20, p. 13], and the fact, noted already, that $\underline{\bigoplus_{I}\mathbb{Z}}=\bigoplus_{I}\underline{\mathbb{Z}}$ for any set $I$. More particularly, letting $\underline{\mathcal{P}}$, much as in the proof of Lemma 3.2, denote the condensation of a free resolution of $\underline{A}$, we have:
@@
and this chain of equalities, if sound, will complete the argument. Let us check the details. Observe first that $\underline{\mathcal{P}}$ is a projective resolution of $\underline{A}$, by [13, Prop. 2.18] and Proposition 12.3 below; this shows the third equality. The second is the aforementioned adjunction. For the first, we have noted already that $\mathbb{Z}[\underline{S}]$ is flat, so the question reduces to that of whether the terms of $\underline{\mathcal{P}}\otimes\mathbb{Z}[\underline{S}]$ are projective. For this observe simply that $\big{(}\underline{\bigoplus_{I}\mathbb{Z}}\big{)}\otimes\mathbb{Z}[\underline{S}]=\big{(}\bigoplus_{I}\underline{\mathbb{Z}}\big{)}\otimes\mathbb{Z}[\underline{S}]=\bigoplus_{I}\big{(}\underline{\mathbb{Z}}\otimes\mathbb{Z}[\underline{S}]\big{)}=\bigoplus_{I}\mathbb{Z}[\underline{S}]$ for any set $I$; that such sums are projective follows from Proposition 12.3 below. ∎
Note that the lemma allows us to recast this paper’s essential background in uniformly classical terms: Shelah’s result was that the implication
@@
is indeed a $\mathsf{ZFC}$ theorem. This implication is, moreover, straightforward to prove, since by a result of Nöbeling ([16], see also [6, Proposition XI.4.4]), generalizing work of Specker [23], the group $C(S,\mathbb{Z})$ is itself a free abelian group, one isomorphic to the free abelian group on the topological weight of $S$.
@@
Lemma 11.1 also sheds additional light on the results of Section 10, both providing an alternative proof of a variation on the main result of that section, as well as showing that the existence of large cardinals provides a bound on the extent to which this result can be generalized.
@@
On the other hand, the existence of sufficiently large cardinals provides a limit on the extent to which the phenomena of Corollary 11.3 can occur. In particular, if $\kappa$ is a strongly compact cardinal, then $\mathsf{CondAb}_{\kappa}$ is sufficiently rich to guarantee a positive answer to the appropriate interpretation of Whitehead’s problem. The following corollary yields clause (3) of Theorem 1.4.
@@
Suppose that $\underline{A}$ is Whitehead in $\mathsf{CondAb}_{\kappa}$. In particular, for for every $S\in\mathsf{ED}_{\kappa}$, we have $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})(S)=0$. By Lemma 11.1, it follows that, for every cardinal $\lambda<\kappa$, we have $\mathrm{Ext}^{1}_{\mathsf{Ab}}(A,\mathbb{Z}^{(\lambda)})=0$. By the observations in the paragraph immediately preceding Corollary 11.2, it follows that $A$ is $\kappa$-free, i.e., every subgroup of $A$ of cardinality ${<}\kappa$ is free. By [6, Corollary II.3.11], since $\kappa$ is strongly compact, $A$ is free, as desired. [^5] As [6, Corollary II.3.11] makes clear, we could have merely assumed that $\kappa$ is only $L_{\omega_{1}\omega}$-compact, a weaker but lesser-known hypothesis than the strong compactness of $\kappa$.
@@
In the proof of Corollary 11.4, the crucial property of the cardinal $\kappa$ was the fact that every $\kappa$-free group is free. This is true if $\kappa$ is a strongly compact cardinal, but we remark that, consistently, this can also be true for smaller values of $\kappa$. In particular, it is shown in [12, §4] that, assuming the consistency of the existence of infinitely many supercompact cardinals, there exists a model of $\sf ZFC$ in which this condition holds for the first cardinal fixed point of the $\aleph$ function, i.e., the least $\kappa$ such that $\kappa=\aleph_{\kappa}$.
12. Locally compact abelian groups
@@
By the structure theorem for locally compact abelian groups (see [20, Thm. 4.1(i)]), any locally compact abelian group $A$ is of the form $\mathbb{R}^{n}\times B$ for some $n\in\mathbb{N}$, where $B$ is an extension of a compact group by a discrete one, or where, in other words, there exists a short exact sequence
@@
with $C$ compact and $D$ discrete whose image in $\mathsf{CondAb}$ is exact as well. [^6] See [13, Props. 4.9, 2.18]. Since $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{\mathbb{R}},\underline{\mathbb{Z}})=0$ [20, p. 25], we have
$(\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})=0)\Leftrightarrow(\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{B},\underline{\mathbb{Z}})=0).$
Consider the following portion of the long exact sequence deriving from equation (1):
@@
(The rightmost 00 follows from [20, p. 27] (e.g., from the comment following Cor. 4.9); the leftmost 00 follows from [20, Prop. 4.2].) Hence
@@
By Theorem 1.2, item (1) holds if and only if $D=\bigoplus_{I}\mathbb{Z}$ for some $I$. Also, the argument of [13, Prop. 4.11] together with [20, Prop. 4.3(ii)] shows that $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{C},\underline{\mathbb{R}})=0$; this together with the fact that $\underline{\mathrm{Hom}}(\underline{C},\underline{\mathbb{R}})=\underline{\mathrm{Hom}(C,\mathbb{R})}=0$ ([20, Prop. 4.2]) implies that the map $\underline{\mathrm{Hom}}(\underline{C},\underline{\mathbb{R}/\mathbb{Z}})\to\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{C},\underline{\mathbb{Z}})$ appearing in the long exact sequence induced by applying $\underline{\mathrm{Hom}}(\underline{C},-)$ to $\underline{\mathbb{Z}}\to\underline{\mathbb{R}}\to\underline{\mathbb{R}/\mathbb{Z}}$ is an isomorphism. Again applying [20, Prop. 4.2], we have $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{C},\underline{\mathbb{Z}})=\underline{\mathrm{Hom}(C,\mathbb{R}/\mathbb{Z})}=\underline{\mathbb{D}(C)}$, the condensed image of the Pontryagin dual group of $C$; in particular, item (2) holds if and only if $C=0$. Summing up, $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})=0$ if and only if $A$ is of the form $\mathbb{R}^{n}\times D$ for some $D=\bigoplus_{I}\mathbb{Z}$. As noted, by [15, Prop. 3.3], such $A$ are precisely the projective objects of $\mathsf{LCA}$. ∎
@@
To see this, recall that, as noted, the compact projective condensed abelian groups of the form $\mathbb{Z}[\underline{S}]$ with $S\in\mathsf{ED}$ generate $\mathsf{CondAb}$ ([3, p. 16–17 and Proposition 3.5]); in particular, for any condensed abelian group $X$ there exists a set $I$, a collection $\{S_{i}\mid i\in I\}$ of spaces in $\mathsf{ED}$, and an epimorphism $r:\bigoplus_{I}\mathbb{Z}[\underline{S_{i}}]\to X$. If $X$ is projective then $r$ admits a right inverse $s$, hence:

### v3

@@ 10. Martin’s Axiom and $\mathsf{CondAb}_{\omega_{1}}$
For a given abelian group $A$, if $\underline{A}$ is Whitehead in $\mathsf{CondAb}_{\omega_{1}}$, then $A$ is Whitehead in the classical sense, since
@@
Let $\vec{p}=\langle p_{\eta}\mid\eta<\omega_{1}\rangle$ be a sequence of conditions in $\mathbb{P}$. By (the proof of) [9, Lemma 7.5], by thinning out $\vec{p}$ to some cofinal subsequence if necessary, we can assume that there is a free subgroup $A^{\prime}$ of $A$ that is pure in $A$ and such that $A_{p_{\eta}}\subseteq A^{\prime}$ for all $\eta<\omega_{1}$ (this is the only use of the full power of Chase’s condition in the proof). Let $B^{\prime}$ be a basis for $A^{\prime}$. Using Claim 10.5 and extending each $p_{\eta}$ if necessary, we can assume that, for all $\eta<\omega_{1}$, $A_{p_{\eta}}$ is freely generated by some finite subset $B^{\prime}_{\eta}$ of $B^{\prime}$. By thinning out $\vec{p}$ again if necessary, we can assume that the sets $\{B^{\prime}_{\eta}\mid\eta<\omega_{1}\}$ form a $\Delta$-system, with root $R$.
@@ 11. A different vantage point
To see the fifth equality, apply Lemma 3.2; for the third and fourth, respectively, apply [18, Prop. 3.5 and Prop. 3.2]. The first equality follows from [26, p. 13 (iii) (cf. p. 25)]. This leaves us only with the second. Here the idea, evidently, is to apply the Hom-tensor adjunction at the level of $\mathrm{Ext}^{1}_{\mathsf{CondAb}}$, and we may do so by the naturality of the adjunction at the $\mathrm{Hom}_{\mathsf{CondAb}}$-level, the flatness of $\mathbb{Z}[\underline{S}]$ [26, p. 13], and the fact, noted already, that $\underline{\bigoplus_{I}\mathbb{Z}}=\bigoplus_{I}\underline{\mathbb{Z}}$ for any set $I$. More particularly, letting $\underline{\mathcal{P}}$, much as in the proof of Lemma 3.2, denote the condensation of a free resolution of $\underline{A}$, we have:
@@
and this chain of equalities, if sound, will complete the argument. Let us check the details. Observe first that $\underline{\mathcal{P}}$ is a projective resolution of $\underline{A}$, by [18, Prop. 2.18] and Proposition 12.3 below; this shows the third equality. The second is the aforementioned adjunction. For the first, we have noted already that $\mathbb{Z}[\underline{S}]$ is flat, so the question reduces to that of whether the terms of $\underline{\mathcal{P}}\otimes\mathbb{Z}[\underline{S}]$ are projective. For this observe simply that $\big{(}\underline{\bigoplus_{I}\mathbb{Z}}\big{)}\otimes\mathbb{Z}[\underline{S}]=\big{(}\bigoplus_{I}\underline{\mathbb{Z}}\big{)}\otimes\mathbb{Z}[\underline{S}]=\bigoplus_{I}\big{(}\underline{\mathbb{Z}}\otimes\mathbb{Z}[\underline{S}]\big{)}=\bigoplus_{I}\mathbb{Z}[\underline{S}]$ for any set $I$; that such sums are projective follows from Proposition 12.3 below. ∎
Note that the lemma allows us to recast this paper’s shaping background in uniformly classical terms: Shelah’s result was that the implication
@@
is indeed a $\mathsf{ZFC}$ theorem. This implication is, moreover, straightforward to prove, since by a result of Nöbeling ([22], see also [10, Proposition XI.4.4]), generalizing work of Specker [29], the group $C(S,\mathbb{Z})$ is itself a free abelian group, one isomorphic to the free abelian group on the topological weight of $S$.
@@
Lemma 11.1 also sheds additional light on the results of Section 10, entailing both an alternative proof of a variation of that section’s main result and limits, in the presence of sufficiently large cardinals, on the extent to which this result can be generalized.
@@
On the other hand, the existence of sufficiently large cardinals will bound the extent to which the phenomena of Corollary 11.3 can occur. In particular, if $\kappa$ is a strongly compact cardinal, then $\mathsf{CondAb}_{\kappa}$ is sufficiently rich to guarantee a positive answer to the appropriate interpretation of Whitehead’s problem. The following corollary yields clause (3) of Theorem 1.4.
@@
Suppose that $\underline{A}$ is Whitehead in $\mathsf{CondAb}_{\kappa}$. In particular, for for every $S\in\mathsf{ED}_{\kappa}$, we have $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})(S)=0$. By Lemma 11.1, it follows that, for every cardinal $\lambda<\kappa$, we have $\mathrm{Ext}^{1}_{\mathsf{Ab}}(A,\mathbb{Z}^{(\lambda)})=0$. By the observations in the paragraph immediately preceding Corollary 11.2, it follows that $A$ is $\kappa$-free, i.e., every subgroup of $A$ of cardinality ${<}\kappa$ is free. By [10, Corollary II.3.11], since $\kappa$ is strongly compact, $A$ is free, as desired. [^5] As [10, Corollary II.3.11] makes clear, we could have merely assumed that $\kappa$ is only $L_{\omega_{1}\omega}$-compact, a weaker but lesser-known hypothesis than the strong compactness of $\kappa$.
@@
In the proof of Corollary 11.4, the crucial property of the cardinal $\kappa$ was the fact that every $\kappa$-free group is free. This is true if $\kappa$ is a strongly compact cardinal, but we remark that, consistently, this can also be true for smaller values of $\kappa$. In particular, it is shown in [16, §4] that, assuming the consistency of the existence of infinitely many supercompact cardinals, there exists a model of $\sf ZFC$ in which this condition holds for the first cardinal fixed point of the $\aleph$ function, i.e., the least $\kappa$ such that $\kappa=\aleph_{\kappa}$.
12. Locally compact and solid abelian groups
@@
By the structure theorem for locally compact abelian groups (see [26, Thm. 4.1(i)]), any locally compact abelian group $A$ is of the form $\mathbb{R}^{n}\times B$ for some nonnegative integer $n$, where $B$ is an extension of a compact group by a discrete one, or where, in other words, there exists a short exact sequence
@@
with $C$ compact and $D$ discrete whose image in $\mathsf{CondAb}$ is exact as well (see [18, Props. 4.9, 2.18]). Since $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{\mathbb{R}},\underline{\mathbb{Z}})=0$ [26, p. 25], we have
$\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})=0\;\Leftrightarrow\;\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{B},\underline{\mathbb{Z}})=0.$
Consider next the following portion of the long exact sequence deriving from equation (1):
@@
(The rightmost 00 follows from [26, p. 27] (e.g., from the comment following Cor. 4.9); the leftmost 00 follows from [26, Prop. 4.2].) Hence
@@
By Theorem 1.2, item (1) holds if and only if $D=\bigoplus_{I}\mathbb{Z}$ for some $I$. Also, the argument of [18, Prop. 4.11] together with [26, Prop. 4.3(ii)] shows that $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{C},\underline{\mathbb{R}})=0$; this together with the fact that $\underline{\mathrm{Hom}}(\underline{C},\underline{\mathbb{R}})=\underline{\mathrm{Hom}(C,\mathbb{R})}=0$ ([26, Prop. 4.2]) implies that the map $\underline{\mathrm{Hom}}(\underline{C},\underline{\mathbb{R}/\mathbb{Z}})\to\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{C},\underline{\mathbb{Z}})$ appearing in the long exact sequence induced by applying $\underline{\mathrm{Hom}}(\underline{C},-)$ to $\underline{\mathbb{Z}}\to\underline{\mathbb{R}}\to\underline{\mathbb{R}/\mathbb{Z}}$ is an isomorphism. Again applying [26, Prop. 4.2], we have $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{C},\underline{\mathbb{Z}})=\underline{\mathrm{Hom}(C,\mathbb{R}/\mathbb{Z})}=\underline{\mathbb{D}(C)}$, the condensed image of the Pontryagin dual group of $C$; in particular, item (2) holds if and only if $C=0$. Summing up, $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})=0$ if and only if $A$ is of the form $\mathbb{R}^{n}\times D$ for some $D=\bigoplus_{I}\mathbb{Z}$. As noted, by [21, Prop. 3.3], such $A$ are precisely the projective objects of $\mathsf{LCA}$. ∎
@@
To see this, recall that, as noted, the compact projective condensed abelian groups of the form $\mathbb{Z}[\underline{S}]$ with $S\in\mathsf{ED}$ generate $\mathsf{CondAb}$ ([7, p. 16–17 and Proposition 3.5]); in particular, for any condensed abelian group $X$ there exists a set $I$, a collection $\{S_{i}\mid i\in I\}$ of spaces in $\mathsf{ED}$, and an epimorphism $r:\bigoplus_{I}\mathbb{Z}[\underline{S_{i}}]\to X$. If $X$ is projective then $r$ admits a right inverse $s$, hence:
@@
Theorem 12.1 opens onto broader questions about alternative interpretations of Whitehead’s problem in the condensed setting. These may be thought of as defined by two main parameters: the subscript $s$ of the $\underline{\mathrm{Ext}}^{1}$ term, and the domain $d$ of $\underline{\mathrm{Ext}}^{1}_{s}(\cdot,\underline{\mathbb{Z}})$. After all, in retrospect, it is not altogether surprising that ($s=\mathsf{CondAb}$)-valued invariants $\underline{\mathrm{Ext}}^{1}_{s}$ should be strong enough to settle the problem over the comparatively weaker domains of $d=\mathsf{Ab}$ or even $d=\mathsf{LCA}$ (we hope to have shown that it is altogether interesting nevertheless). A natural next question, then, is how much further this $d$ may be pushed. By Corollary 12.2, of course, Whitehead’s problem has a ZFC solution when $d=\mathsf{CondAb}$ as well, but it is an unsatisfying one, deriving as it does from the arguably spurious case of $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{\mathbb{R}},\underline{\mathbb{Z}})=0$. Against this background, particularly interesting is the problem when $d$ equals the subcategory $\mathsf{Solid}$ of $\mathsf{CondAb}$ mentioned briefly in our introduction. Forming, as it does, the paradigmatic example of an analytic ring, this is one of $\mathsf{CondAb}$’s most critical subcategories (see [26, §5–7] and [25, §2]); moreover, as noted, it was this category’s features which Clausen and Scholze leveraged in their original proof of Theorem 1.2. The solid setting, being “very much a ‘nonarchimedean’ notion” [26, footnote 10 on p. 32], is also one in which $\underline{\mathbb{R}}$ does not appear. [^6] More precisely, the inclusion $\mathsf{Solid}\to\mathsf{CondAb}$ possesses a left adjoint solidification functor notated $M\mapsto M^{\blacksquare}$, and the solidification $\underline{\mathbb{R}}^{\blacksquare}$ of $\underline{\mathbb{R}}$ is 00 [26, Corollary 6.1(iii)]. Within it, lastly, $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}$ and $\underline{\mathrm{Ext}}^{1}_{\mathsf{Solid}}$ coincide [26, Lemma 5.9], so that this instance of the problem recovers the $s=d$ symmetry of the classical one:
Question 12.5.
We thank Peter Scholze for answering with a multifaceted no; the following observations are essentially his.
First, for the solid abelian group $A=\prod_{\mathbb{N}}\bigoplus_{\mathbb{N}}\underline{\mathbb{Z}}$ and any $n\geq 0$,
$\underline{\mathrm{Ext}}^{n}_{\mathsf{CondAb}}(A,\underline{\mathbb{Z}})(S)\cong\mathrm{lim}^{n}\,\mathbf{A}[C(S,\mathbb{Z})],$
where $\mathbf{A}[H]$ denotes for any abelian group $H$ the inverse system indexed by functions $f:\mathbb{N}\to\mathbb{N}$, with terms $A_{f}:=\bigoplus_{i\in\mathbb{N}}H^{(f(i))}$ and natural projection maps $A_{g}\to A_{f}$ for any $f$ which is pointwise dominated by $g$. See [5] for details of this conversion; the immediate point, though, is that these derived limits have been the subject of much recent study (see [4, 3, 33]), from which the consistency with the ZFC axioms of the statement that
$\mathrm{lim}^{1}\,\mathbf{A}[H]=0\;\text{ but }\mathrm{lim}^{2}\,\mathbf{A}[H]\neq 0$
for every nontrivial abelian group $H$, for example, is known. In fact this holds in any model of the Proper Forcing Axiom; whenever it does, $A$ will clearly witness a negative answer to Question 12.5.
Left open in the preceding analysis is the possibility of a consistently affirmative answer to Question 12.5, since
$\mathrm{lim}^{n}\,\mathbf{A}[H]=0\text{ for all }n>0\text{ and abelian groups }H$
and hence
$\underline{\mathrm{Ext}}^{n}_{\mathsf{Solid}}(A,\underline{\mathbb{Z}})=0\text{ for all }n>0$
is also consistent with the ZFC axioms [1, Thm. 3]. By the following argument, though, $A$ witnesses a negative answer to Question 12.5 in this scenario as well.
Consider the surjection
$t:\bigoplus_{\mathbb{N}^{\mathbb{N}}}\prod_{\mathbb{N}}\mathbb{Z}\to\prod_{\mathbb{N}}\bigoplus_{\mathbb{N}}\mathbb{Z}$
defined as follows: fix $g:\mathbb{N}\to\mathbb{N}$ and an element $a$ of the $g^{\mathrm{th}}$ summand of the domain of $t$; clearly it will suffice to describe the effect of $t$ on such $a$. Here it is convenient to adopt the (elsewhere innocuous) convention that 00 is not a natural number: the first $g(1)$ many entries in $a$ are then mapped to the first $g(1)$ summands of the first factor of $\prod_{\mathbb{N}}\bigoplus_{\mathbb{N}}\mathbb{Z}$, the next $g(2)$ many entries to the first $g(2)$ many summands of the second factor of $\prod_{\mathbb{N}}\bigoplus_{\mathbb{N}}\mathbb{Z}$, and so on. Equipping $\mathbb{Z}$ with the discrete topology and letting products and sums induce the product and coproduct topologies, respectively, we may view $t$ as a continuous mapping of topological abelian groups. Consider the condensed image $\underline{t}$ of $t$: observe that its codomain is $A$, its domain is solid, and that it remains an epimorphism (cf. [18, Prop. 2.18]). Thus, to show that $A$ is not projective in $\mathsf{Solid}$, it will suffice to show that $\underline{t}$ does not split, and for this it will suffice to show that the mapping $\underline{t}(*)=t$ of discrete abelian groups described above possesses no right-inverse $s$. By [10, Cor. III.1.9], though, for any
$s:\prod_{\mathbb{N}}\bigoplus_{\mathbb{N}}\mathbb{Z}\to\bigoplus_{\mathbb{N}^{\mathbb{N}}}\prod_{\mathbb{N}}\mathbb{Z}$
there exists an $m\in\mathbb{N}$ and finite $F\subseteq\mathbb{N}^{\mathbb{N}}$ such that the image of the restriction of $s$ to $\prod_{n\geq m}\bigoplus_{\mathbb{N}}\mathbb{Z}$ is contained within $\bigoplus_{f\in F}\prod_{\mathbb{N}}\mathbb{Z}$. Letting $k=\max\{f(m)\mid f\in F\}+1$ and writing $x$ for any element of $\prod_{n\geq m}\bigoplus_{\mathbb{N}}\mathbb{Z}$ which is nonzero at the $k^{\mathrm{th}}$ summand of the $m^{\mathrm{th}}$ factor, we must then have $ts(x)\neq x$. This shows that $t$ possesses no right-inverse, [^7] One can alternatively rule out such a right-inverse via the $\alpha=2$ instance of [10, Thm. X.2.3]. and hence that $A$ is not projective, completing the argument.
Even this doesn’t altogether settle the question, though, since it is also consistent with the ZFC axioms (and follows from the continuum hypothesis) that $\underline{\mathrm{Ext}}^{1}(A,\underline{\mathbb{Z}})\neq 0$ [17, Thm. 2]. For a model-independent negative answer to Question 12.5, take the solid abelian group $B=\underline{\mathbb{Z}_{p}}/\underline{\mathbb{Z}[1/q]}$ for any distinct prime numbers $p$ and $q$. As noted in Clausen’s Lecture 11 of [8],
$\mathrm{Ext}^{2}_{\mathsf{CondAb}}(B,\underline{\mathbb{Z}})=\mathbb{Z}_{q}/\mathbb{Z}[1/p],$
hence $B$ is not projective; it remains only to check that it is Whitehead. In fact, for the numerator
[…]
of $B$, we have
[…]
$=(\underline{\mathbb{Z}[1/p]/\mathbb{Z}})[-1]$
(see [19, Rmk. 2, p. 2] for the first equality), and for its denominator
[…]
we have
$R\underline{\mathrm{Hom}}_{\mathsf{CondAb}}(\underline{\mathbb{Z}[1/q]},\underline{\mathbb{Z}})$ $=R\mathrm{lim}\;(R\underline{\mathrm{Hom}}_{\mathsf{CondAb}}(\underline{\mathbb{Z}},\underline{\mathbb{Z}})\xleftarrow{\times q}R\underline{\mathrm{Hom}}_{\mathsf{CondAb}}(\underline{\mathbb{Z}},\underline{\mathbb{Z}})\leftarrow\cdots)$
$=R\mathrm{lim}\;(\underline{\mathbb{Z}}\xleftarrow{\times q}\underline{\mathbb{Z}}\xleftarrow{\times q}\cdots)$
$=(\underline{\mathbb{Z}_{q}/\mathbb{Z}})[-1].$
Hence, applying $R\underline{\mathrm{Hom}}_{\mathsf{CondAb}}(\,\cdot\,,\underline{\mathbb{Z}})$ to the short exact sequence
$0\to\underline{\mathbb{Z}[1/q]}\to\underline{\mathbb{Z}_{p}}\to\underline{\mathbb{Z}_{p}}/\underline{\mathbb{Z}[1/q]}\to 0$
we find that
$R\underline{\mathrm{Hom}}_{\mathsf{CondAb}}(\underline{\mathbb{Z}_{p}}/\underline{\mathbb{Z}[1/q]},\underline{\mathbb{Z}})=(\underline{\mathbb{Z}_{q}}/\underline{\mathbb{Z}[1/p]})[-2]$
and that $\underline{\mathrm{Ext}}^{1}(B,\underline{\mathbb{Z}})$, in particular, equals zero.